OpenStudy (anonymous):

How to solve this differential equation: 2 dy/dx = 1-y*2

OpenStudy (anonymous):

Yes.

OpenStudy (unklerhaukus):

\[2 \frac{\text dy}{\text dx} = 1-y^2\] \[ \frac{2\text dy}{1-y^2} = \text dx\]

OpenStudy (turingtest):

most likely...

OpenStudy (unklerhaukus):

i dont get it

OpenStudy (turingtest):

partial fractions?

OpenStudy (unklerhaukus):

break up the square?

OpenStudy (turingtest):

\[\frac{2}{1-y^2}=\frac A{1+y}+\frac B{1-y}\]yeah

OpenStudy (anonymous):

I don't know how to do it.

OpenStudy (anonymous):

Can someone please show me how to answer this question.

OpenStudy (unklerhaukus):

\[\frac{2}{1-y^2}=\frac A{1+y}+\frac B{1-y}\] \[2=A({1-y})+B({1+y})\] \[2=(A+B)+y(B-A)\]

OpenStudy (unklerhaukus):

\[A+B=2\]\[B-A=0\]

OpenStudy (unklerhaukus):

A=B=1

OpenStudy (anonymous):

I know how to do the partial fraction part but I'm stuck after that.

OpenStudy (unklerhaukus):

\[\frac{2}{1-y^2}=\frac 1{1+y}+\frac 1{1-y}\] \[\int\frac 1{1+y}+\frac 1{1-y}\text dy = \int\text dx\]

OpenStudy (unklerhaukus):

\[\int \frac{\text dy}{1+y}+-\int\frac{-\text dy}{1-y}=\int x\]

OpenStudy (unklerhaukus):

\[\ln|1+x|-\ln|1-y|=x+c\]

OpenStudy (unklerhaukus):

*\(\int dx\)

OpenStudy (unklerhaukus):

\[\ln\left|\frac{1+x}{1-y}\right|=x+c\] \[\frac{1+y}{1-y}=e^ce^x\] \[\frac{1+y}{1-y}=Ae^x\]

OpenStudy (unklerhaukus):

*\(\ln|1+y|\)

OpenStudy (anonymous):

In your third step, why did you bring the minus sign out. why the (+) sign can't remain ?

OpenStudy (turingtest):

\[u=1-y\implies du=-dy\]he just wrote it in such a way as to not make you do the -substitutiuon

OpenStudy (turingtest):

u-substitution*

OpenStudy (unklerhaukus):

it is a strange step , and i dont really get it my self, if you try to integrate with out the negative you get arc-tans instead of logs

OpenStudy (anonymous):

Yes, that's what I mean. If the (+) sign remains then the answer will be totally different, Anyway thanks.

OpenStudy (unklerhaukus):

i think it might have something to do with the domain of x